Thank you so much for your help...
Select the values of x and y that make this equation true.
6:27:12=24:x:y

Again, thank you so much, Rebecca

Answer by JBarnum(2146)

6*4=24
x=4*27=108
y=4*12=48
6:27:12=24:108:48